# When Black Holes Relax in a Cold Bath: Evolving Page Curves for Black Holes Coupled to Cooling Baths

### v1

@@ I Introduction
In Almheiri:2019hni , for example, the set-up considered was a two-dimensional black hole in Jackiw-Teitelboim gravity attached to a bath. By computing the quantum extremal surface via the generalized entropy formula of Engelhardt and WallEngelhardt:2014gca , it was found that there are two extremal surfaces that correspond to two saddles of the gravitational path integral—the Hawking saddle and the island saddle. A switch between the saddles takes place at Page time, and the Page curve is recovered. Since these breakthroughs, a deeper understanding of the transition has been obtained in terms of replica wormholes Penington:2019kki ; Almheiri:2019qdq ; Almheiri:2020cfm . Building on these developments, there have been many investigations of Page curves for various black holes Almheiri:2019yqk ; Marolf:2020xie ; Almheiri:2019psy ; Chen:2020uac ; Hashimoto:2020cas ; Hartman:2020khs ; Gautason:2020tmk ; Balasubramanian:2020xqf ; Krishnan:2020fer ; Krishnan:2020oun ; Ling:2020laa ; Ahn:2021chg .
Generally, the evaporation of a black hole will be influenced by external conditions, such as ambient temperature or energy exchange with other systems. It is interesting to ask how changing the external conditions affects the Page curve? In this paper, we address this question for black holes in JT gravity, where we can model the changing external conditions by changing the bath temperature.
We consider that the bath cools down, its temperature continuously evolving with time via some given rule. As the temperature of the bath changes, it leads to a change in the boundary parameterization $f(t)$ that describes the black hole geometry in JT gravity. This in turn modifies the positions of the quantum extremal surface (QES) at different times and hence the Page curve. We compute the modified Page curves numerically.
@@
In previous literature, a scenario in a somewhat similar spirit was explored in Hollowood:2020cou , where the modification of the Page curve due to shock waves sent from the bath to the black hole were considered. In this paper, which obtained analytical results, the increase of temperature was assumed to be instantaneous. Our investigations are numerical, which allows us to go beyond these limiting assumptions and investigate the most general cases where the temperature of the bath is changed continuously with time.
@@
II Preliminaries
II.1 Jackiw-Teitelboim Gravity
Here, we outline the basic aspects of Jackiw-Teitelboim gravity(JT gravity). We refer the reader to Mertens:2022irh for a recent detailed review.
@@ II.1 Jackiw-Teitelboim Gravity
II.2 Quantum Extremal Surfaces
@@
II.3 Black Hole at Equilibrium
II.3.1 Geometry
In this section, we briefly introduce the setup. In our case, we have the AdS2 Poincaré patch with coordinates $x^{\pm}$:
@@ II.3.1 Geometry
An eternal black hole with temperature $\frac{1}{\beta}$ has the boundary parametrization relating the Poincaré and black hole patches given by Hollowood:2020cou .
@@
II.3.2 Stress Tensors
The ADM energy of spacetime is holographically related to the boundary parameterizationMaldacena:2016upp
@@ II.4 Non-Equilibrium
II.5 Early Times and Late Times
@@
III Evaporation of black hole: instantaneous change in bath temperature
@@ III.2 Island saddle
IV Black Hole Evaporation: General Cooling Profiles
@@
IV.1 Bath Energy follows Newton’s law of cooling
@@
Figure 8(a)8(b)8(c) corresponds to Energy, Hawking saddle entropy, and Island saddle entropy, respectively. We see that the island entropy curve is slightly concave at early times and nearly flat at late times.
@@
‘Figure 9(a)9(b)9(c) corresponds to Energy, Hawking saddle entropy, and Island saddle entropy, respectively. We can see that the island entropy curve is nearly flat at early times and convex at late times.
@@
Figure 10(a)10(b)10(c) corresponds to Energy, Hawking saddle entropy, and Island saddle entropy, respectively. We find that the island entropy curve is slightly concave at early times and slightly convex at late times.
IV.2 Linear Decrease in Incoming Stress Tensor.
@@
Figure 11(a)11(b)11(c) corresponds to Energy, Hawking saddle entropy, and Island saddle entropy, respectively. We find the entropy curve to be slightly concave throughout. We do not expect the curve to be convex at later times, as there is no final temperature to settle in.
IV.3 Quasi-static Evaporation

### v3

# Kashmir-190006, India

@@ 1 Introduction
In Almheiri:2019hni , for example, the set-up considered was a two-dimensional black hole in Jackiw-Teitelboim gravity attached to a bath. By computing the quantum extremal surface via the generalized entropy formula of Engelhardt and WallEngelhardt:2014gca , it was found that there are two extremal surfaces corresponding to two saddles of the gravitational path integral—the Hawking saddle and the island saddle. A switch between the saddles takes place at Page time, and the Page curve is recovered. Since these breakthroughs, a deeper understanding of the transition has been obtained in terms of replica wormholes Penington:2019kki ; Almheiri:2019qdq ; Almheiri:2020cfm . Building on these developments, there have been many investigations of Page curves for various black holes Almheiri:2019yqk ; Marolf:2020xie ; Almheiri:2019psy ; Chen:2020uac ; Hashimoto:2020cas ; Hartman:2020khs ; Gautason:2020tmk ; Balasubramanian:2020xqf ; Krishnan:2020fer ; Krishnan:2020oun ; Ling:2020laa ; Ahn:2021chg .
In general, the evaporation of a black hole will be influenced by external conditions, such as ambient temperature or energy exchange with other systems. It is interesting to ask how changing the external conditions affects the Page curve? In this paper, we address this question in the same set-up involving a black hole coupled to a bath used in Almheiri:2019hni , where we model the changing external conditions by changing the bath temperature.
We consider a bath whose temperature continuously evolves with time via some given formula. As the temperature of the bath changes, it leads to a change in the boundary parameterization $f(t)$ that describes the black hole geometry in JT gravity. This in turn modifies the positions of the quantum extremal surface (QES) at different times, and hence the Page curve is also modified. We compute the modified Page curves numerically.
@@
A scenario in a somewhat similar spirit was explored in Hollowood:2020cou . This paper considered the modification of the Page curve due to shock waves sent from the bath to the black hole. In this paper, which obtained analytical results, the increase of temperature was assumed to be instantaneous. Our investigations are numerical, which allows us to go beyond these limiting assumptions and investigate more general cases where the temperature of the bath is changed continuously with time.
@@
2 Preliminaries
2.1 Jackiw-Teitelboim Gravity
Here, we outline the basic aspects of Jackiw-Teitelboim(JT) gravityJackiw:1984je ; Teitelboim:1983ux and nearly AdS2 holographyMaldacena:2016hyu ; Almheiri:2014cka ; Jensen:2016pah ; Maldacena:2016upp ; Engelsoy:2016xyb . We refer the reader to Mertens:2022irh for a recent detailed review.
@@ 2.1 Jackiw-Teitelboim Gravity
2.2 Quantum Extremal Surfaces
@@
2.3 Black Hole at Equilibrium
We now introduce the setup of a black hole coupled to a bath. The framework of JT gravity black holes with leaky boundary conditions was originally developed in Engelsoy:2016xyb ; Mertens:2019bvy .
In our case, we have the AdS2 Poincaré patch with coordinates $x^{\pm}$:
@@ 2.3 Black Hole at Equilibrium
An eternal black hole with temperature $\frac{1}{\beta}$ has the boundary parametrization relating the Poincaré and black hole patches given by:
@@
The ADM energy of spacetime is holographically related to the boundary parameterization:
@@ 2.4 Non-Equilibrium
2.5 Early Times and Late Times
@@
3 Evaporation of black hole: instantaneous change in bath temperature
@@ 3.2 Island saddle
4 Black Hole Evaporation: General Cooling Profiles
@@
4.1 Bath Energy follows Newton’s law of cooling
@@
Figure 8 shows how the energy, Hawking saddle entropy, and Island saddle entropy varies with time in the $k>s$ case. We see that the island entropy curve is slightly concave at early times and nearly flat at late times.
@@
Figure 9 shows how the relevant quantities vary with time when $k<s$. We can see that the island entropy curve is nearly flat at early times and convex at late times.
@@
The behavior of the relevant quantites for $k>s$ case is summarised in Figure 10. We find that the island entropy curve is slightly concave at early times and slightly convex at late times.
4.2 Linear Decrease in Incoming Stress Tensor.
@@
Figure 11 shows the change in energy, Hawking saddle entropy, and Island saddle entropy with time for this case. We find the entropy curve to be slightly concave throughout. We do not expect the curve to be convex at later times, as there is no final temperature to settle in.
4.3 Quasi-static Evaporation